# Optimal trajectories for Bayesian olfactory search in turbulent flows: the low information limit and beyond

### v2

@@ Abstract
For robotics applications, this observation may be used to formulate biomimetic strategies, as in [14]. Alternatively, if the searcher is equipped with a model for the spatially-dependent likelihood of an encounter, the search problem may be formalized as a Bayesian partially-observable Markov decision process (POMDP) [19]. This formulation allows the optimal, i.e., minimum average time, search strategy to be described by a dynamic programming equation, which may be approximately solved [20, 21, 22].
@@
where $\theta$ is a Heaviside function, we have supposed the source is in $x=0$, $y$ is the distance from the plume centerline, $\alpha\geq 0$ is a free parameter that depends on the environment, and $x_{D}$ is a characteristic decay length which depends on $c_{\rm thr}$. This result is only valid sufficiently far from the source.
@@
Rather than impose one of these models directly, we used empirical models taken from simulations. We performed high-fidelity direct numerical simulations (DNS) of a turbulent flow (${\rm Re}_{\lambda}\simeq 150$), while tracking the positions of tracer particles which are emitted by a stationary point source. The system is advected by a uniform mean flow $U\hat{x};$ we repeated the simulation for $U/u_{\rm rms}\in\{0,1.2,2.5,4.9,7.4\}$. The tracers are taken as proxies for a passive scalar cue; this is a good approximation at high Péclet number. After coarse-graining the particle data onto a 2–D grid of about $\simeq 3250$ cells, we imposed a threshold concentration $c_{\rm thr}=200$ particles/cell and computed the likelihood $p_{1}(\bm{x})$ that this threshold is instantaneously exceeded (see Fig. 1). For the largest values of $U,$ the data are a good fit to Eq. 2 with $\alpha=1,$ and for the smallest values of $U$ the data fit Eq. 3 (see Supplemental Material).
@@
For each flow, we then used the SARSOP algorithm [26] to extract a quasi-optimal search policy from the underlying dynamic programming equation [27, 19] (see Supplemental Material). To study trajectories, we generated large ensembles of trajectories via Monte Carlo trials; for comparison, we did the same for the popular heuristic policy “infotaxis” [28], which seeks to minimize the entropy of the posterior. During search, observations were randomly drawn from the empirical likelihood. This artificially suppressed correlations between observations, but we found that correlations do not strongly affect optimal trajectories qualitatively; instead, their main effect is on the typical arrival time to the source. Therefore, we defer our study of the effect of correlations to a forthcoming paper, where searches will be performed in the DNS data.
@@
Results. Our results can be organized and understood through the following principle: because encounters are rare, the most important optimization the searcher must perform is to decide how it should move when it does not smell anything, and the single most likely trajectory is the one generated in the absence of any encounters after time zero. We will call this deterministic trajectory the “no-hit trajectory.” To support this claim, we evaluated the probability $P_{\rm nh}(t)$ that, at time $t,$ the searcher has not encountered the cue since time 0, given that it has not yet found the source. This is plotted for the largest wind speed in the upper panel of Fig. 3; the results for other wind speeds are very similar and shown in the Supplemental Material. For both quasi-optimal and infotactic motion, $P_{\rm nh}(t)\gtrsim 0.5$ for all times $t$, so the no-hit trajectory plays an outsize role.
Figure 3: Search trajectory statistics for $U/u_{\rm rms}\simeq 7.4.$ Top panel: probability $P_{\rm nh}(t)$ that, at time $t,$ the searcher has not encountered the cue since time 0, given that it has not yet found the source. Lower panels: mean squared displacement (MSD) in the $x$ and $y$ directions.
Figure 4: Quasi-optimal no-hit trajectories, for $U/u_{\rm rms}\simeq 7.4$ The searcher is shown as a red square, the searcher’s belief is shown in grayscale, and the starting point is a blue circle.
@@
Other examples of no-hit trajectories are supplied in the Supplemental Material. As the wind speed becomes smaller, the downwind return occurs earlier and earlier, and zigzagging quickly becomes irrelevant to the optimal motion. Instead, at moderate wind speed, a quasi-optimal searcher exhibits repeated up- and downwind motion that begins to resemble a complex, asymmetric spiral. We suggest that the curves tend to approximately lie on isolines of the prior, which would imply the trajectories are essentially performing gradient ascent of the likelihood profile. When $U$ and $u_{\rm rms}$ are comparable, the problem has very little symmetry and, accordingly, the trajectories are especially complex. At $U=0,$ both the quasi-optimal and infotaxis trajectories resemble Archimedean spirals, consistent with previous studies [28, 31, 20].
@@
Low-information limit. The results at large wind speed can be understood by means of a simplified model. In particular, we again assume the searcher initializes its belief as the prior $b_{0}(\bm{x})$ induced by the encounter at time zero, i.e., proportional to Eq. 2 for large enough $x.$ However, consistent with the dominance of the no-hit trajectory, we assume successive encounters only occur when the searcher is close to the source—the likelihood is very close to unity in a strip directly downwind of the source (see yellow region in Fig. 1, lower panel), which is where $\langle c\rangle>c_{\rm thr}.$ Empirically (and unsurprisingly), this strip has negligible length except when $U$ is large.
@@
where $x$ and $y$ are, respectively the upwind and crosswind position of the source with respect to the searcher, and $\ell^{*}$ is the characteristic length of the strip. The searcher uses this likelihood to update its belief. We will call this simplified model the low-information limit; A similar model was used to describe sector search in Ref. [32].
In this limit, the result of the Bayesian update (Eq. 1) is that the searcher will simply zero out its prior in a $\ell^{*}$-neighborhood (and renormalize) as it moves (see upper panel of Fig. 5). The optimal trajectory will then be the one that most efficiently explores the prior’s support, so the geometry of the prior plays an essential role.
@@
We now determine the optimal trajectories in the low-information limit. Due to the Gaussian factor in Eq. 2, the prior will be strongly confined to a cone $C$ given by $|y|<ax$ with $a\sim u_{\rm rms}/U,$ so it is useful to approximate the confinement as exact: suppose the prior depends only on $x$ inside the cone, and it is zero outside (see upper panel of Fig. 5). We seek to minimize the expected time of arrival
@@
A zigzag trajectory can be (for large enough $x$) modeled by a sequence of purely crosswind moves at positions $x_{1},x_{2},\dots$ (see Fig. 5, lower left panel). A key observation is that the searcher should not move further upwind than $\ell^{*}$ during any crosswind traversal, i.e., we have the constraint $x_{k+1}-x_{k}\leq\ell^{*}$ for each $k$. Otherwise, there will be points not swept out by the trajectory, and because the trajectory is monotonic in $x,$ $J$ will be infinite.
In the absence of the inequality constraint, the decision problem of choosing the optimal sequence $\{x_{k}\}$ is exactly equivalent to the linear search problem (LSP), a well-studied decision problem which asks for the turning points of a trajectory on a line which minimizes the expected arrival time to a target drawn from some known probability density [34, 35, 36, 37]. In the Supplemental Material we prove that if $p(x)$ decays exponentially or slower—as it does in our case—then $x_{k+1}-x_{k}$ asymptotically diverges for large $k.$ Thus the inequality constraint must saturate far enough upwind, i.e., $x_{k+1}-x_{k}=\ell^{*},$ and one can show this gives rise to diffusive scaling as observed in the MSD data. The resulting trajectory may be modeled as (ignoring numerical factors)
@@
with asymptotically constant speed $\sim v$ in an rms sense (here, $v$ is again the speed of the searcher).
We therefore conclude that the zigzag motion with $t^{1/2}$ scaling is a consequence of the fact that (a) the likelihood decays exponentially with $x$ (and not faster), (b) the likelihood decays faster than exponential outside the cone, and (c) the likelihood is small outside of a strip in front of the source. However, the argument (in particular, point (b)) rests on the approximation $\exp[-(u_{\rm rms}y/Ux)^{2}]\ll\exp(-x/x_{D}).$ When the searcher reaches an upwind distance $x_{R}\sim x_{D},$ this approximation breaks down, and the local prior probability will be comparable to that just outside the cone, near the starting point (see Fig. 5, lower right panel). This means that there is no continued benefit to search upwind, and the searcher should instead return downwind—as we indeed observe in the quasi-optimal policy.
We cannot analytically predict the optimal trajectories at finite $U$ (and moreover the approximation Eq. 4 becomes untenable), but the low-information limit still grants some insights. Using the analysis of Ref. [15], one expects $x_{D}\sim u_{\rm rms}^{-1}(SU/c_{\rm thr})^{1/2}$ (where $S$ is the emission rate) when $U$ is large enough that ballistic single- and two-particle dispersion scaling still holds on relevant timescales. This is in basic agreement with our observation that downwind returns occur earlier as $U$ decreases. Eventually, as $U$ becomes comparable to $u_{\rm rms},$ diffusive motion begins to dominate trajectories of Lagrangian particles on timescales of interest. In this regime, the likelihood is no longer well-confined to a cone and gains some support even upwind of the source (this is visible in our simulations at $U/u_{\rm rms}\simeq 1.2$). This complexification of the prior geometry leads to concomitantly complex optimal trajectories (see the Supplemental Material).

### v3

@@ Abstract
For robotics applications, this observation may be used to formulate biomimetic strategies, as in [14]. Alternatively, if the searcher is equipped with a model for the spatially-dependent likelihood of an encounter, the search problem may be formalized as a Bayesian partially-observable Markov decision process (POMDP) [19]. This formulation allows the optimal, i.e., minimum average time, search strategy to be described by a dynamic programming (Bellman) equation, which may be approximately solved [20, 21, 22].
@@
where $u_{\rm rms}$ is the typical fluctuation speed of the flow, $\theta$ is a Heaviside function, we have supposed the source is in $x=0$, $y$ is the distance from the plume centerline, $\alpha\geq 0$ is a free parameter that depends on the environment, and $x_{D}$ is a characteristic decay length which depends on $c_{\rm thr}$. This result is only valid sufficiently far from the source.
@@
Rather than impose one of these models directly, we used empirical models taken from simulations. We performed high-fidelity direct numerical simulations (DNS) of a turbulent flow (${\rm Re}_{\lambda}\simeq 150$), while tracking the positions of tracer particles which are emitted by a stationary point source. The system is advected by a uniform mean flow $U\hat{x};$ we repeated the simulation for $U/u_{\rm rms}\in\{0,1.2,2.5,4.9,7.4\}$. The tracers are taken as proxies for a passive scalar cue; this is a good approximation at high Péclet number. After coarse-graining the particle data onto a 2–D grid of about $\simeq 3250$ cells, we imposed a threshold concentration $c_{\rm thr}=200$ particles/cell and computed the likelihood $p_{1}(\bm{x})$ that this threshold is instantaneously exceeded (see Fig. 1). For the largest values of $U,$ the data are a good fit to Eq. 2 with $\alpha=1,$ and for the smallest values of $U$ the data fit Eq. 3 (see Fig. 3 of the Supplemental Material  [^1] See Supplemental Material at [hyperlink], which includes Refs. [43, 44], for detailed information about methods for the DNS and POMDP solution, as well as additional calculations, plots, and movies.).
@@
For each flow, we then used the SARSOP algorithm [27] to extract a quasi-optimal search policy from the underlying dynamic programming equation [28, 19] (Supplemental Material, Eq. 3 [26]). To study trajectories, we generated large ensembles of trajectories via Monte Carlo trials; for comparison, we did the same for the popular heuristic policy “infotaxis” [29], which seeks to minimize the entropy of the posterior.
@@
Results. Our results can be organized and understood through the following principle: because encounters are rare, the most important optimization the searcher must perform is to decide how it should move when it does not smell anything, and the single most likely trajectory is the one generated in the absence of any encounters after time zero. We will call this deterministic trajectory the “no-hit trajectory.” To support this claim, we evaluated the probability $P_{\rm nh}(t)$ that, at time $t,$ the searcher has not encountered the cue since time 0, given that it has not yet found the source. This is plotted for the largest wind speed in the upper panel of Fig. 4; the results for other wind speeds are very similar and shown in Figs. 4–5 in the Supplemental Material [26]. For both quasi-optimal and infotactic motion, $P_{\rm nh}(t)\gtrsim 0.5$ for all times $t$, so the no-hit trajectory plays an outsize role.
Figure 3: Quasi-optimal no-hit trajectory, for $U/u_{\rm rms}\simeq 7.4$ The searcher is shown as a red square, the searcher’s belief is shown in grayscale, and the starting point is a blue circle.
Figure 4:
@@
Other examples of no-hit trajectories are supplied in Secs. VI C–D of the Supplemental Material [26]. As the wind speed decreases, the downwind return occurs earlier and earlier, and zigzagging quickly becomes irrelevant to the optimal motion. Instead, at moderate wind speed, a quasi-optimal searcher exhibits repeated up- and downwind motion that begins to resemble a complex, asymmetric spiral. We suggest that the curves tend to approximately lie on isolines of the prior, which would imply the trajectories are essentially performing gradient ascent of the likelihood profile. When $U$ and $u_{\rm rms}$ are comparable, the problem has very little symmetry and, accordingly, the trajectories are especially complex. At $U=0,$ both the quasi-optimal and infotaxis trajectories resemble Archimedean spirals, consistent with previous studies [29, 32, 20].
@@
Low-information limit. The results at large wind speed can be understood by means of a simplified model. In particular, we again assume the searcher initializes its belief as the prior $b_{0}(\bm{x})$ induced by the encounter at time zero, i.e., proportional to Eq. 2 for large enough $x.$ However, consistent with the dominance of the no-hit trajectory, we assume successive encounters only occur when the searcher is close to the source—the likelihood is very close to unity in a strip directly downwind of the source (see yellow region in Fig. 1, lower panel), which is where $\langle c\rangle>c_{\rm thr}.$ Empirically, this strip has negligible length except when $U$ is large;
@@
where $x$ and $y$ are, respectively the upwind and crosswind position of the source with respect to the searcher, and $\ell^{*}$ is the characteristic length of the strip. The searcher uses this likelihood to update its belief. We will call this simplified model the low-information limit. A similar model was used to describe sector search in Ref. [33].
In this limit, the result of the Bayesian update (Eq. 1) is that the searcher will simply zero out its prior in a $\ell^{*}$-neighborhood (and renormalize) as it moves (see Fig. 5, panel (a)). The optimal trajectory will then be the one that most efficiently explores the prior’s support, so the geometry of the prior plays an essential role.
@@
We now determine the optimal trajectories in the low-information limit. Due to the Gaussian factor in Eq. 2, the prior will be strongly confined to a cone $C$ given by $|y|<ax$ with $a\sim u_{\rm rms}/U,$ so it is useful to approximate the confinement as exact: suppose the prior depends only on $x$ inside the cone, and it is zero outside (see Fig. 5, panel (a)). We seek to minimize the expected time of arrival
@@
A zigzag trajectory can be (for large enough $x$) modeled by a sequence of purely crosswind moves at positions $x_{1},x_{2},\dots$ (see Fig. 5, panel (b)). A key observation is that the searcher should not move further upwind than $\ell^{*}$ during any crosswind traversal, i.e., we have the constraint $x_{k+1}-x_{k}\leq\ell^{*}$ for each $k$. Otherwise, there will be points not swept out by the trajectory, and because the trajectory is monotonic in $x,$ $J$ will be infinite.
In the absence of the inequality constraint, the decision problem of choosing the optimal sequence $\{x_{k}\}$ is exactly equivalent to the linear search problem (LSP), a well-studied decision problem which asks for the turning points of a trajectory on a line which minimizes the expected arrival time to a target drawn from some known probability density [35, 36, 37, 38]. This is shown in the Supplemental Material (Sec. V), after which we prove that if $p(x)$ decays exponentially or slower—as it does in our case—then $x_{k+1}-x_{k}$ asymptotically diverges for large $k$ [26]. Thus the inequality constraint must saturate far enough upwind, i.e., $x_{k+1}-x_{k}=\ell^{*},$ and one can show this gives rise to diffusive scaling as observed in the MSD data.
(ignoring numerical factors)
@@
with asymptotically constant speed $\sim v$ in an rms sense.
We therefore conclude that the zigzag motion with $t^{1/2}$ scaling is a consequence of the fact that (a) the likelihood decays exponentially with $x$ (and not faster), (b) the likelihood decays faster than exponential outside the cone, and (c) the likelihood is small outside of a strip in front of the source. However, the argument (in particular, point (b)) rests on the approximation $\exp[-(u_{\rm rms}y/Ux)^{2}]\ll\exp(-x/x_{D}).$ When the searcher reaches an upwind distance $x_{R}\sim x_{D},$ this approximation breaks down, and the local prior probability will be comparable to that just outside the cone, near the starting point (see Fig. 5, panel (c)). This means that there is no continued benefit to search upwind, and the searcher should instead return downwind—as we indeed observe in the quasi-optimal policy.
We cannot analytically predict the optimal trajectories at finite $U$ (and moreover the approximation Eq. 4 becomes untenable), but the low-information limit still grants some insights. Using the analysis of Ref. [15], one expects $x_{D}\sim u_{\rm rms}^{-1}(SU/c_{\rm thr})^{1/2}$ (where $S$ is the emission rate) when $U$ is large enough that ballistic single- and two-particle dispersion scaling still holds on relevant timescales. This is in basic agreement with our observation that downwind returns occur earlier as $U$ decreases. Eventually, as $U$ becomes comparable to $u_{\rm rms},$ diffusive motion begins to dominate trajectories of Lagrangian particles on timescales of interest. In this regime, the likelihood is no longer well-confined to a cone and gains some support even upwind of the source (this is visible in our simulations at $U/u_{\rm rms}\simeq 1.2$). This complexification of the prior geometry leads to concomitantly complex optimal trajectories (see Supplemental Material, Sec. VI A, C, D [26]).